4. (3 pts) A soft drink manufacturer currently makes the aluminum cans for its beverages. It has the following costs for maki this is a management systems control subject
Answer #1

Cost of buying 2,000,000 cans=$0.18*2000000=$360,000 or $360K

If company buy the cans, variable costs will be saved, fixed costs will incur irrespective whether company buys or makes.

Savings in variable cost=170+50+100=$320K

We find that savings in variable cost is lower than the cost of buying the cans. So, it is better to make the cans.